OpenStudy (anonymous):

Solve the following logarithmic equation. log₃(x+4) = 2+log₃(x-20)

OpenStudy (anonymous):

\[\log_{3}(x+4)=2+\log_{3}(x-20) \]

OpenStudy (anonymous):

bring the log on the right over to the left side and rewrite as 1 log... \(\large log_{3}(x+4)=2+\log_{3}(x-20)\) \(\large log_{3}(x+4)-\log_{3}(x-20)=2\) \(\large log_{3}(\frac{x+4}{x-20})=2\) \(\large \frac{x+4}{x-20}=3^2 \) can you take it from here?